Circular No. 6935

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

MICROLENSING EVENT A. Becker, on behalf of the MACHO Collaboration (cf. IAUC 6845),

reports the discovery of a likely exotic gravitational microlensing event in progress toward the Small Magellanic Cloud: “The star in question is located at α = 0^{h}45^{m}35^{s}.2, δ = −72^{o}52′34′′ (equinox 2000.0). The star was constant at R = 21.7, V = 21.8 during 1993–1998 and was observed to have brightened significantly on 1998 May 25. Follow-up observations taken with the Cerro Tololo 0.9-m telescope indicate a strong deviation from standard microlensing occuring near June 5, where the star brightened from a magnification of 3 to > 13. This is thought to have been a caustic crossing due to a binary lensing system. Preliminary binary microlensing fits by the MACHO team and S. Rhie indicate that a second caustic crossing will occur within ∼ 2 weeks, possibly between June 21 and 27, although the date is extremely uncertain. This will consist of a sharp rise and fall in brightness of > 2 mag within a few hours. Very frequent photometry (hourly or more often) during the second caustic crossing will provide valuable information on the transverse velocity of the lensing system, which helps to constrain the lens mass and distance. Spectroscopy of the brightened star will also be important for estimating the source angular diameter. A finding chart and additional information are available by anonymous ftp

at darkstar.astro.washington.edu:/macho/Alerts/98-SMC-1 or at the

MACHO Web site (cf. IAUC 6312, 6720). Observers are asked to contact macho@astro.washington.edu, A. Becker at telephone 206-543-9205, or D. Bennett at telephone 219-631-8298, to coordinate observations.”

1998 KY_{26}

S. Ostro, L. Benner, J. Giorgini, K. Rosema, and D. Yeomans, Jet Propulsion Laboratory, write: “We have detected Goldstone (8510-MHz, 3.5-cm) radar echoes from 1998 KY_{26} (MPEC 1998-L02) on June 6–8 that have a bandwidth > 11 Hz, providing the constraint P < 0.55D, where P is rotation period in min and D is the diameter in m (Ostro et al. 1995, Science 270, 80). Delay-Doppler images with a range resolution of 19 m do not resolve the asteroid, suggesting a diameter no greater than 40 m and therefore P < 22 min.”

1998 June 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green